# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

### v3

@@ Abstract
As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it for $r\in\mathbb{N}$ by means of M-L bounds on integrals as
$|A_{r}|\leq c^{r}$
@@
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
@@
${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
and neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. Then
@@
is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
@@
The binary and unary operations on $R[[x]]$ listed so far, ++, $\cdot$, $(\cdot)^{\prime}$, and $\int(\cdot)$, are formal in the sense that, for example, for every $n\in\mathbb{N}_{0}$ the coefficient
$x^{n}](f\cdot g)$
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other operations. In formal proofs involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$, only formal operations may be used, and it is forbidden to use the functional values
@@
However, we can present two such proofs in the next two sections if formality is relaxed to semiformality. Semiformal operations (in Section 3 we introduce several of them) are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$. Semiformal proofs concerning elements in $K[[x]]$ are proofs that satisfy the following condition.
Only formal and semiformal operations on $K[[x]]$ are used. For every formal power series involved in the proof, only at most countably many of its values are used.
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental. We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j\in[t]$ ($\{1,2,\dots,t\}$), are $2t$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
@@ 3 The proof by formal integration
We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [7]. The idea is to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce further operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we mentioned in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
The operation $f(x)\mapsto f(ax)$, which we call the constant inner product, is formal. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
@@
If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
@@
It is easy to show that $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and that $\int\mathrm{Exp}=\mathrm{Exp}-1$.
We introduce a semiformal operation.
Definition 3.3 (semiformal shifts)
@@
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and semiformal shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
In the next two propositions, we show that semiformal shifts behave in desirable ways.
@@
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=}$
$=[x^{n}]f(x+(c+d))\,.$
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$. Thus
@@
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=}$
$=[x^{n}](fg)(x+c)\,.$
@@
The following semiformal exponential identity is somewhat counter-intuitive (at first we got it wrong).
@@
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
@@
We introduce Newton integrals of convergent formal power series.
@@
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. The Newton integral is a semiformal operation because it uses two values of a formal power series. We prove three integral transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
@@
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we add
@@
In the fifth line, we use the binomial theorem. $\Box$
We introduce improper Newton integrals of convergent formal power series.
@@
We leave the proof of the independence of the limit $L$ on the sequences $(b_{n})$ to the reader as an exercise.
@@
We need to bound Newton integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)$ is in $\mathbb{Z}[x]$. We easily obtain the next bound based on two values of the primitive.
@@
It is crucial that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ is Euler’s identity. We prove it in the language of improper Newton integrals of convergent formal power series.
@@
We are ready to translate Hilbert’s proof of the transcendence of $\mathrm{e}$ into formal power series. We again assume for the contradiction that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as at the beginning, only now they are defined in an AB &\& SF way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
@@
Using Propositions 3.12 and 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
@@
The previous proof is semiformal and not formal due to the use of various semiformal operations and many values of formal power series. It is clearly an AB proof.

### v4

@@ Abstract
As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it by means of M-L bounds on integrals as
$|A_{r}|\leq c^{r},\ r\in\mathbb{N}\,$,
@@
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. We may work with functionals on $R[[x]]$, which are maps from $R[[x]]$ to $R$. For example, for every $n\in\mathbb{N}_{0}$ we have the functional $f\mapsto[x^{n}]f$.
Besides functionals, the set $R[[x]]$ is endowed with the operations of addition and multiplication
@@
${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}}$
which have neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. For any $c,d\in R$ and $f,g\in R[[x]]$ we may form the linear combination
$cf(x)+dg(x)=\sum_{n\geq 0}(ca_{n}+db_{n})x^{n}\,.$}
This is in fact a family of binary operations on $R[[x]]$, one for every pair $c,d$. The structure
@@
is a domain, in which $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
@@
The functionals and operations on $R[[x]]$ we have listed so far, $f\mapsto[x^{n}]f$, ++, $\cdot$, $(\cdot)^{\prime}$, $\int(\cdot)$ and linear combinations, are formal in the sense that, for example, every coefficient
$x^{n}](f\cdot g),\ n\in\mathbb{N}_{0}\,$,
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other functionals and operations. We say that a proof involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ is formal if it employs only formal functionals and operations, and if it does not use any functional values
@@
However, we can present two such proofs in the next two sections if formality is relaxed to semiformality. Semiformal functionals are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$. The same holds for semiformal operations. For examples of semiformal functionals and operations on $\mathbb{R}[[x]]$ see Definitions 3.7 and 3.3, respectively. We say that a proof involving elements in $K[[x]]$ is semiformal if it satisfies the following condition.
Only formal and semiformal functionals and operations on $K[[x]]$ are used in the proof. For every formal power series involved in the proof, only at most countably many of its values are used.
@@ 2 Specializing the proof of Beukers, Bézivin and Robba
We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental. We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j=1,2,\dots,t$, are $2t\geq 2$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
@@ 3 The proof by formal integration
We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [7]. The idea is to translate Hilbert’s proof, described in Section 1, into the language of formal power series.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce further functionals and operations on $\mathbb{R}[[x]]$ besides those mentioned in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
@@
The family of operations
$\{f(x)\mapsto f(ax)\colon\;a\in\mathbb{R}\}\,$,
which we call constant inner products, consists of formal unary operations. For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
@@
Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
@@
Clearly, $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and $\int\mathrm{Exp}=\mathrm{Exp}-1$.
We introduce a family of semiformal unary operations.
Definition 3.3 (shifts)
@@
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
We next show that shifts behave in desirable ways.
@@
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=[x^{n}]f(x+(c+d))\,.}$
@@
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. Then the coefficients $x^{n}]f(x+c)$ and $x^{n}]g(x+c)$ equal to
$\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,$,}
respectively. Thus
@@
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}}$
${\textstyle=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=[x^{n}](fg)(x+c)\,.}$
@@
The following semiformal exponential identity is perhaps counter-intuitive; at first we got it wrong.
@@
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
@@
We introduce a family of semiformal functionals on $\mathbb{R}[[x]]$.
@@
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. We prove three transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
@@
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we added
@@
In the fifth line, we used the binomial theorem. $\Box$
We introduce another family of semiformal functionals on $\mathbb{R}[[x]]$.
@@
We leave the proof of the independence of the limit $L$ on the sequence $(b_{n})$ to the reader as an exercise.
@@
We need bounds, obtained in an AB way, on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
@@
As we know, it is crucial that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof is Euler’s identity. We prove it now in an AB &\& SF way.
@@
We are ready to state Hilbert’s proof in an AB &\& SF way. We again assume for the contradiction that
@@
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an AB &\& SF way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
@@
Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
@@
The previous proof is semiformal and not formal due to the use of semiformal functionals and operations, and use of many values of formal power series. It is clearly an AB proof.